# Decentralized Reliability Estimation for Mixnets

### v1

@@ Appendix C Approaches to setting $\hat{\beta}_{e}$
C.1 Naive approach: $\hat{\beta}_{e}=1$
@@
Such model is however easily subverted in a strategic adversarial setting: a malicious node $i$ can selectively drop received packets (after storing their tag in the Bloom filter) destined to one or more target successors $j$, in order to unfairly degrade their measured performance $\hat{\rho}_{j}$. Given a mixnet with layer width $W$, an adversary that controls a set of $i\in A$ nodes in a layer can degrade the measured performance $\hat{\rho}_{j}$ of any number of nodes $j$ in the succeeding layer, by up to $\frac{|A|}{W}$ at zero cost for the adversarial nodes in terms of lower measured performance $\hat{\rho}_{i}$ for malicious nodes $i\in A$.
C.2 Symmetric approach: $\hat{\beta}_{e}=\frac{1}{2}$
An alternative approach to penalize and discourage such adaptive malicious behaviour is to define node performance considering that half the ‘blame’ for drops in link $e=(i,j)$ is with $i$, and the other half with $j$, i.e., setting $\hat{\beta}_{e}=\frac{1}{2}$ for all links $e$. This symmetric attribution approach is effective at discouraging malicious packet drops as any drop now affects the performance of the adversarial node comparably to that of its target, which can be either a predecessor or a successor of the malicious node. We note that the symmetric attribution of drops however leads to an increasingly asymmetric performance penalty for large scale attacks involving many adversarial nodes and many targets. In such cases, $\hat{\beta}_{e}=0.5$ $\forall e\in G$ begins to impose a larger performance penalty on link successors than it does on predecessors. This is due to successors receiving significantly less total traffic (due to the large scale nature of the attack) and thus the attributed half of the drops represent a larger fraction of the total inputs than for the predecessors, to whom the other half of drops are attributed.
A practical concern in cases where no attack takes place is that the symmetric attribution of blame results in a lowered performance $\hat{\rho}_{i}$ for the predecessors $i$ of nodes $j$ that are either offline or congested and unable to process all the received traffic – a scenario that may not be so rare. This may unfairly degrade $\hat{\rho}_{i}<{\rho}_{i}$ for reliable nodes $i$ at the expense of overestimating $\hat{\rho}_{j}>{\rho}_{j}$ for unreliable nodes $j$.
@@ C.2 Symmetric approach: $\hat{\beta}_{e}=\frac{1}{2}$
C.3 Threshold approach: $\hat{\beta}_{e}\in\{0,\frac{1}{2},1\}$
The threshold approach aims to: (i) eliminate the unfair degradation of measured performance for reliable nodes due to successors being offline or congested, or predecessors being consistently faulty; while (ii) still symmetrically penalizing malicious nodes that conduct selective dropping attacks to degrade the estimated performance of honest, reliable predecessors or successors. This is achieved by assigning values to $\hat{\beta}_{e}$ that take into account node performance across its whole set of incoming or outgoing edges. The performance $\hat{\rho}_{e}$ of each edge is computed with Eq. 4.
C.3.1 Median incoming and outgoing performance
@@
C.3.2 Threshold criteria
Let $\bar{\tau}$ be a threshold for the median performance that nodes are expected to achieve at a minimum when participating in the mixnet. The $\hat{\beta}_{e}$ is set per edge $e=(i,j)$ and it takes values in $\{0,\frac{1}{2},1\}$, according to the following criteria:
  • $\hat{\beta}_{(i,j)}=1$: if $\bar{\rho}^{\mathsf{in}}_{j}<\bar{\tau}$ and $\bar{\rho}^{\mathsf{out}}_{i}\geq\bar{\tau}$ ($j$ has sub-par median incoming performance $\bar{\rho}^{\mathsf{in}}_{j}$ while $i$ has above-threshold outgoing median performance $\bar{\rho}^{\mathsf{out}}_{i}$). If $\sum_{i}{s}_{(i,j)}^{*}=0$ then $\hat{\beta}_{(i,j)}=1$ for any value of $\bar{\rho}^{\mathsf{out}}_{i}$.
  • $\hat{\beta}_{(i,j)}=0$: if $\bar{\rho}^{\mathsf{in}}_{j}\geq\bar{\tau}$ and $\bar{\rho}^{\mathsf{out}}_{i}<\bar{\tau}$ ($j$ has above-threshold median incoming performance while $i$ is showing low performance in more than half of its outgoing links).
  • $\hat{\beta}_{(i,j)}=\frac{1}{2}$: if $\bar{\rho}^{\mathsf{out}}_{i},\bar{\rho}^{\mathsf{in}}_{j}\geq\bar{\tau}$ (both $i$ and $j$ have above-threshold median performance), or $\bar{\rho}^{\mathsf{out}}_{i},\bar{\rho}^{\mathsf{in}}_{j}<\bar{\tau}$ (both $i$ and $j$ have sub-par median performance).
@@ C.3.2 Threshold criteria
Assuming that in every layer half the nodes are reliable, a reliable node $j$ has median incoming performance $\bar{\rho}^{in}_{j}\geq\bar{\tau}$, as well as median outgoing performance $\bar{\rho}^{\mathsf{out}}_{j}\geq\bar{\tau}$. In this situation, any adversarial predecessor or successor that also has above-threshold median performance is penalized symmetrically with $\hat{\beta}_{e}=\frac{1}{2}$, taking upon itself as much blame for dropped packets as it inflicts on the target, which acts as a disincentive for the attack. If the adversarial node is instead underperforming in the median due to dropping too many packets, then the adversary receives all the blame for the drops, without affecting the performance of the reliable honest targets.
The $\bar{\tau}$ threshold also bounds the performance hit taken by reliable honest nodes due to non-malicious random failures of other nodes. One scenario is when a node $j$ goes offline during the epoch, and thus fails to receive packets from all predecessors $i\in P(j)$ while offline. The threshold approach avoids penalizing all the well-performing predecessors $i$ (that have $\bar{\rho}^{\mathsf{out}}_{i}\geq\bar{\tau}$) for these dropped packets by setting $\hat{\beta}_{(i,j)}=1$ for those links. Another likely scenario for non-malicious failure is congestion, i.e., a node $j$ lacking the throughput to process all the packets routed through it, and thus failing to receive and process a non-negligible fraction of the packets sent by predecessors $i$. Due to widespread packet losses by $j$, we have again that $\bar{\rho}^{\mathsf{in}}_{j}<\bar{\tau}$, and thus all the packet drops due to node $j$’s limited capacity are attributed to $j$ rather than to its well-performing predecessors $i$, by setting $\hat{\beta}_{(i,j)}=1$. Similarly, a node $i$ with consistent faulty behaviour in its outgoing links $(i,j)$, i.e., with $\bar{\rho}^{\mathsf{out}}_{i}<\bar{\tau}$ and a pattern of recording received packets in its Bloom filter but somehow consistently failing to successfully get those packets delivered to the next hop $j$, is also blamed for the drops $\hat{d}_{(i,j)}$, by setting $\hat{\beta}_{(i,j)}=0$ on $i$’s outgoing links, so that reliable successors $j$ do not take any performance hit from $i$’s consistently faulty behaviour.
C.3.3 Choice of threshold $\bar{\tau}$
@@
The choice of value for $\bar{\tau}$ has the following effects. First, assuming that at least half of each layer is honest and reliable, the measured performance for a reliable node $j$ is always $\hat{\rho}_{j}\geq\bar{\tau}$. Therefore, a higher $\bar{\tau}$ more tightly bounds the unfair blame that perfect nodes may receive in cases where multiple predecessors or successors perform just above the threshold $\bar{\tau}$. However, if the threshold performance $\bar{\tau}$ is too high to be attainable in practice, then many nodes will be in the ‘underperforming’ category. For nodes within this group, the performance hit due to others’ drops is potentially much larger, since underperforming nodes are blamed by either half or all of the drops in their links (depending on whether the node at the other end of the link is also underperforming). In the extreme situation where all nodes fail to reach $\bar{\tau}$, the scheme behaves as in the symmetric approach. Thus, minimizing the impact on the measured performance for perfect nodes by setting a very high $\bar{\tau}$ comes at the expense of more likely and potentially larger performance impact for slightly-less-than-perfect nodes. In our experimental evaluations we use $\bar{\tau}=0.99$ as threshold.
Appendix D Mixnet Simulator
@@
This section provides a description of the simulator we implemented [^8] The code can be made available upon request to reviewers and we plan to submit it for artifact evaluation. and used to conduct an empirical analysis of the proposed node performance estimation protocol. The scripts are in Python and use the SimPy library to handle events such as packet creation, packet sending, receiving and dropping. This section provides extensive details on the features, parameters and experimental setups we used in to obtain the results presented in previous sections, as well as additional experiments excluded from the main body due to page limits.
D.1 Clients and Encoded Packets
The simulator runs clients that encode and send packets throughout a simulated epoch of one hour. Packet generation by clients follows a Poisson process with a configurable rate of $\lambda$ packets per second. Each encoded packet is a measurement packet with (globally configurable) probability $p_{\mathsf{lot}}$. The rest are data packets addressed to other (randomly chosen) clients or cover traffic that clients send to themselves.
@@ D.1 Clients and Encoded Packets
Packets routes are selected according to a uniform routing policy, considering a mixnet with $L=3$ layers and $W=80$ mix nodes per layer, as well as $W_{G}=80$ gateways. [^9] To make the simulations as realistic as possible, we select network and mixing delay parameters similar to those in the deployed Nym network. We consider that legitimate client traffic is uniformly distributed over all gateways and also addressed in equal measure to recipients in all gateways. The encoded delay per packet per mix node is exponentially distributed with average $50$ms, and we additionally consider $40$ms of transmission time in each link and $2$ms of packet processing time at the gateways.
D.2 Reliable, Unreliable and Malicious Nodes
@@
In the simulations we treat gateways and mix nodes equally in terms of reliability, failure modes and malicious behaviour, and thus refer to all $80$ gateways and $240$ mix nodes indistinctly as ‘nodes’ that receive packets from their predecessors and forward them to their successors, sometimes dropping packets if they are adversarial or not fully reliable. Packets may be dropped by a node before or after the packet tag is stored in the node’s Bloom filter. If a packet is dropped before storing, it is considered that the drop happened in the node’s incoming link. If the drop happens after storing then it is accounted for in the node’s outgoing link.
Nodes can have various types of behaviours:
  • Reliable nodes are honest and function perfectly, meaning that they are always online and correctly process all the received packets without dropping any.
  • Unreliable or faulty nodes are non-adversarial but subject to failures that cause packet drops. The possible failures include: (a) going offline for some (exponentially distributed) period of time, dropping any packets that were inside the node at the moment of going offline, as well as failing to receive new packets while offline; (b) becoming congested due to having a limited throughput and dropping packets on incoming links when the incoming packet rate exceeds the node’s throughput; (c) suffer from random glitches that cause packet drops in incoming or outgoing links with a certain probability.
  • Malicious nodes are always online and have high throughput but they purposefully drop all or a fraction of packets in links with selected predecessor or successor targets, in order to degrade the performance estimates $\hat{\rho}_{j}$ for those targets.
D.3 Simulation Outputs
@@
Once the epoch finalizes, the simulator logs in .csv files the outputs of the sampling protocol, including the ${s}^{*}_{e}$, ${d}^{*}_{e}$ counts per link $e$ as well as $\hat{\rho}_{j}$, $\bar{\rho}^{\mathsf{in}}_{j}$, $\bar{\rho}^{\mathsf{out}}_{j}$ estimates for every node $j$. In addition, the simulator logs packet counts for all edges and nodes that would not be available in an actual deployment – but that are available in a simulated environment and useful for evaluation purposes. This includes ${s}_{e}$, ${d}_{e}$, ${\rho}_{e}$, and ${\beta}_{e}$ for all edges as well as ${\rho}_{j}$ for all nodes and gateways. We evaluate the proposed performance estimation protocols by comparing the $\hat{\rho}_{j}$ output by the protocol to the underlying ground truth performance ${\rho}_{j}$, both in settings where packet drops are due to random failures and adversarial settings where packet drops are malicious.
D.4 Simulation Runtime
@@
D.5 Experimental Setup: Honest but Unreliable Setting
@@
  • $32$ unreliable nodes per layer ($40\%$ of the layer) may go offline during the epoch. Each of these nodes toggles between being online and offline for exponentially distributed lengths of time, with average $90$ minutes online and $10$ minutes offline; i.e., on average, these nodes are online $90\%$ of the time. Within a one-hour epoch however, about half of these nodes stay online the entire epoch and in practice behave as reliable, with the other half going offline for at least part of the epoch.
  • $4$ unreliable nodes per layer ($5\%$ of the layer) have limited throughput, with one of them having a throughput equal to the average per-node incoming packet rate (and thus only dropping some packets when traffic is higher than average), and the other three having throughput that is $\frac{1}{2}$, $\frac{1}{4}$, and $\frac{1}{8}$ of the average incoming packet rate (and thus dropping, respectively, about $\frac{1}{2}$, $\frac{3}{4}$, and $\frac{7}{8}$ of the incoming packets).
  • The final $4$ unreliable nodes per layer ($5\%$ of the layer) randomly drop some packets. One of these nodes drops $1\%$ of the incoming packets, another drops $1\%$ of outgoing packets, a third drops $20\%$ of incoming packets and the last drops $20\%$ of outgoing packets.
@@
We set the probability of a packet being a measurement to $1\%$, i.e., $p_{\mathsf{lot}}=0.01$. We vary the total amount of packets encoded by clients to study the node performance estimation accuracy relative to the number of measurement samples taken throughout the epoch. The results are shown in Figure 2.
D.6 Experimental Setup: Adversarial Setting
To simulate adversarial settings we consider scenarios with $A=\{1,2,4,8,16,32,64\}$ adversaries and $T=\{1,2,4,8,16,32,64\}$ targets. We only consider node assignments to layers that are most favorable to the adversary, i.e., where all $A$ adversarial nodes are in layers adjacent to all the $T$ targets, either as predecessors, as successors, or both. In cases where all $A$ nodes are in a single layer, the $T$ nodes may be all in the preceding layer, all in the succeeding layer or, for $T>1$, distributed as $\frac{T}{2}$ in the preceding and $\frac{T}{2}$ in the succeeding layers. In cases where the $A$ nodes are distributed among two layers, we consider that all $T$ targets are in the layer between the adversaries, with $\frac{A}{2}$ as predecessors and the other $\frac{A}{2}$ as successors of the $T$ targets. For $64$ adversaries or targets we only consider scenarios where they are split over two layers (rather than all in one layer).
@@ D.6 Experimental Setup: Adversarial Setting
We consider that the $A$ adversaries maximize the attack impact by dropping all (rather than just some) of the packets in the edges shared with the $T$ targets – while not dropping any packets on edges shared with non-targets. Targets are reliable nodes that do not drop any packets and thus have a true performance of $100\%$. For the remaining ‘vanilla’ nodes, which are non-adversarial and non-targets, we consider two cases: one where they are all reliable (results reported in Section 4.1.2), and another where $30\%$ of the nodes in each layer are potentially unreliable (results reported in Appendix D.7). The second case uses a setup almost identical to the one described in Section D.5, with the only difference being that only $16$ (rather than $32$) unreliable nodes per layer may go offline during the epoch.
In order to evaluate a large number of simulated scenarios within a reasonable time frame, we set $p_{\mathsf{lot}}=1$ in these simulations, meaning that in practice we exclude non-measurement packets from the simulation. This speeds up by $100$x the time needed to obtain experimental results without sacrificing accuracy of evaluation, since both targets and adversaries are considered to be reliable in the absence of attack and their measured performance loss is thus compared to having a reliable performance of $100\%$. Thus, simulating non-measurement packets does not add any useful information to the output while multiplying simulation time by 100. We consider scenarios with a total of $2$ million measurement packets per epoch of one hour.
D.7 Experimental Results: Settings that Combine Adversarial and Unreliable Nodes
Here we examine scenarios where some nodes not involved in the attack are unreliable, as described in the experimental setup in Appendix D.6. Similar to the evaluation in Section 4.1.2, we compute the adversarial cost for $|A|$ malicious nodes as $c_{A}=|A|-\sum_{i\in A}\hat{\rho}_{i}$ and the cost imposed on the $|T|$ targets as $c_{T}=|T|-\sum_{j\in T}\hat{\rho}_{j}$. Each simulation run produces a result sample $(c_{A},c_{T})$ that we represent as a dot in a scatter plot. We represent scenarios with large number of adversaries with a different symbol to better evaluate the impact of high corruption rates.
The results are shown in Figure 4, where we can see that adversary and target penalties are still symmetric in the majority of scenarios. We can see that $c_{A}\approx c_{T}$ in all scenarios below a certain scale. As the attack scale (proportional to $|A|$x$|T|$) increases however, there are increased chances that unreliable nodes affect the measured performance of either adversaries or targets, making the attack cost asymmetric. This happens in particular when the combination of attack-related drops and unreliability affects the median incoming or outgoing link performance of a node, i.e., when more than half of a node’s incoming or outgoing links have sub-par performance for one or another reason. As explained in Appendix C, this makes the node be labeled as unreliable and increases the attribution of drops to that node. This can more easily happen to targets that are under attack from a large number of adversaries in an adjacent layer, as well as to adversaries that attack a large number of targets.
On the left side of Figure 4 above the diagonal we see a set of results for which the adversary succeeds in causing a large performance degradation on targets at no penalty to itself ($c_{A}<<c_{T}$). This degree of adversarial success requires controlling $40\%$ ($32$ out of $80$ nodes) of at least one layer (all those samples are green ‘+’ symbols) but also some luck for the adversary. In various other instances with the same degree of adversarial control we observe that either: (i) the adversary has a penalty $c_{A}$ that is lower than that of the targets’ $c_{T}$, but still proportional to it and non-negligible (row of green ‘+’ symbols over the diagonal for which $c_{A}<c_{T}$); (ii) the adversary’s penalty is symmetric with the target’s (‘+’ symbols on the diagonal for which $c_{A}\approx c_{T}$); or (iii) in a few cases the adversary pays a large penalty $c_{A}>10$ without inflicting any penalty on the targets (‘+’ symbols on the bottom of the plot for which $c_{A}>>c_{T}$).
The considered level of unreliability is not helpful for adversaries that control up to $20\%$ of any layer (purple circle symbols). In these cases, at best the adversary pays a symmetric penalty (results on the diagonal, for which $c_{A}\approx c_{T}$), while in other instances it pays a higher cost than the target (circles below the diagonal for which $c_{A}>c_{T}$). In some instances the adversary pays a large penalty without the attack having an effect on the target (circles on the bottom of the figure for which $c_{A}>>c_{T}$).
@@ D.7 Experimental Results: Settings that Combine Adversarial and Unreliable Nodes
Figure 4: Results for settings with both adversarial and unreliable nodes. In the scatter plot each sample represents the aggregate performance penalties $c_{T}$ for $1\leq|T|\leq 64$ targets ($y$ axis) and $c_{A}$ for $1\leq|A|\leq 64$ adversaries ($x$ axis) in $630$ simulation runs. Purple circles represent scenarios where adversaries control at most $20\%$ ($16$ out of $80$ nodes) of any layer. Green ‘+’ signs correspond to simulations where adversaries make up $40\%$ ($32$ out of $80$ nodes) of at least one layer.

### v2

@@ A.2 Approaches to setting $\hat{\beta}_{e}$
A.2.1 Naive approach: $\hat{\beta}_{e}=1$
@@
Such model is however easily subverted in a strategic adversarial setting: a malicious node $i$ can selectively drop received packets (after storing their tag in the Bloom filter) destined to one or more target successors $j$, in order to unfairly degrade their measured reliability $\hat{\rho}_{j}$. Given a mixnet with layer width $W$, an adversary that controls a set of $i\in A$ nodes in a layer can degrade the measured reliability $\hat{\rho}_{j}$ of any number of nodes $j$ in the succeeding layer, by up to $\frac{|A|}{W}$ at zero cost for the adversarial nodes in terms of lower measured reliability $\hat{\rho}_{i}$ for malicious nodes $i\in A$.
A.2.2 Symmetric approach: $\hat{\beta}_{e}=\frac{1}{2}$
An alternative approach to penalize and discourage such adaptive malicious behaviour is to define node reliability considering that half the ‘blame’ for drops in link $e=(i,j)$ is with $i$, and the other half with $j$, i.e., setting $\hat{\beta}_{e}=\frac{1}{2}$ for all links $e$. This symmetric attribution approach is effective at discouraging malicious packet drops as any drop now affects the reliability of the adversarial node comparably to that of its target, which can be either a predecessor or a successor of the attacker. A practical concern in cases where no attack takes place is that the symmetric attribution of blame results in a lowered reliability $\hat{\rho}_{i}$ for the predecessors $i$ of nodes $j$ that are either offline or congested and unable to process all the received traffic – a scenario that may not be rare. This may unfairly degrade $\hat{\rho}_{i}<{\rho}_{i}$ for reliable nodes $i$ at the expense of overestimating $\hat{\rho}_{j}>{\rho}_{j}$ for unreliable nodes $j$.
Consider a case where $j\in F$ nodes in a layer of width $W$ go offline shortly after processing a few packets, [^5] In the extreme case where a node $j$ has zero measurement packets in its Bloom filter for the epoch, i.e., $\sum_{i}{s}_{(i,j)}^{*}=0$, then the node is marked as having reliability $\hat{\rho}_{j}=0$, and all drops in the edges $(i,j)$ can be safely attributed to $j$, i.e., $\hat{\beta}_{(i,j)}=1$. while all their predecessors $i$ are honest and reliable throughout the entire epoch, i.e., ${\rho}_{i}=1$ for all nodes in the preceding layer. Attributing drops symmetrically causes $\hat{\rho}_{i}$ to be unfairly degraded by $\frac{|F|}{2W}$, while offline nodes $j$ unfairly get credit for processing half the traffic sent their way, and their measured reliability may be as high as $\hat{\rho}_{j}\approx\frac{1}{2}$ while in truth ${\rho}_{j}\approx 0$.
A.2.3 Threshold approach: $\hat{\beta}_{e}\in\{0,\frac{1}{2},1\}$
The threshold approach aims to: (i) eliminate the unfair degradation of measured reliability for reliable nodes due to successors being offline or congested, or predecessors being consistently faulty; while (ii) still symmetrically penalizing malicious nodes that conduct selective dropping attacks to degrade the estimated reliability of honest, reliable predecessors or successors. This is achieved by assigning values to $\hat{\beta}_{e}$ that take into account node reliability across its whole set of incoming or outgoing edges. The reliability $\hat{\rho}_{e}$ of each edge is computed with Eq. 4.
@@ A.2.3 Threshold approach: $\hat{\beta}_{e}\in\{0,\frac{1}{2},1\}$
Threshold criteria. Let $\bar{\tau}$ be a threshold for the median reliability that nodes are expected to achieve at a minimum when participating in the mixnet. The $\hat{\beta}_{e}$ is set per edge $e=(i,j)$ and it takes values in $\{0,\frac{1}{2},1\}$, according to the following criteria:
@@
[…]
@@
Choice of threshold $\bar{\tau}$. The choice of value for $\bar{\tau}$ has the following effects. First, assuming that at least half of each layer is honest and reliable, the measured performance for a reliable node $j$ is always $\hat{\rho}_{j}\geq\bar{\tau}$. Therefore, a higher $\bar{\tau}$ more tightly bounds the unfair blame that perfect nodes may receive in cases where multiple predecessors or successors perform just above the threshold $\bar{\tau}$. However, if $\bar{\tau}=1$ is rarely attainable in practice, then many nodes will be in the ‘underperforming’ category. In the extreme situation where all nodes fail to reach $\bar{\tau}$, the scheme behaves as in the symmetric approach. In our experimental evaluations we use $\bar{\tau}=0.99$ as threshold.
A.3 Mixnet Simulator
@@
This section provides a description of the simulator we implemented [^6] The code can be made available upon request to reviewers and we plan to make it public with the publication. and used to conduct an empirical analysis of the proposed reliability estimation protocol. The scripts are in Python and use the SimPy library to handle events such as packet creation, packet sending, receiving and dropping. This section provides extensive details on the features, parameters and experimental setups used in to obtain the results presented in previous sections, as well as additional experiments excluded from the main body due to page limit.
A.3.1 Clients and Encoded Packets
@@
The simulator runs clients that encode and send packets throughout a simulated epoch of one hour. Packet generation by clients follows a Poisson process with a configurable rate of $\lambda$ packets per second. Each encoded packet is a measurement packet with (globally configurable) probability $p_{\mathsf{lot}}$. The rest are data packets addressed to other (randomly chosen) clients or cover traffic that clients send to themselves. Packets’ routes are selected according to a uniform routing policy, considering a mixnet with $L=3$ layers and $W=80$ mix nodes per layer, as well as $W_{G}=80$ gateways. [^7] To make the simulations as realistic as possible, we select network and mixing delay parameters similar to those in the deployed Nym network. We consider that legitimate client traffic is uniformly distributed over all gateways and also addressed in equal measure to recipients in all gateways. The encoded delay per packet per mix node is exponentially distributed with average $50$ms, and we additionally consider $40$ms of transmission time in each link and $2$ms of packet processing time at the gateways.
A.3.2 Reliable, Unreliable and Malicious Nodes
@@
In the simulations we treat gateways and mix nodes equally in terms of reliability, failure modes and malicious behaviour, and thus refer to all $80$ gateways and $240$ mix nodes indistinctly as ‘nodes’ that receive packets from their predecessors and forward them to their successors, sometimes dropping packets if they are adversarial or not fully reliable. Packets may be dropped by a node before or after the packet tag is stored in the node’s Bloom filter. If a packet is dropped before storing, it is considered that the drop happened in the node’s incoming link. If the drop happens after storing then it is accounted for in the node’s outgoing link. Nodes can have various types of behaviours:
[…]
A.3.3 Simulation Outputs
@@
Once the epoch finalizes, the simulator logs in .csv files the outputs of the sampling protocol, including the ${s}^{*}_{e}$, ${d}^{*}_{e}$ counts per link $e$ as well as $\hat{\rho}_{j}$, $\bar{\rho}^{\mathsf{in}}_{j}$, $\bar{\rho}^{\mathsf{out}}_{j}$ estimates for every node $j$. In addition, the simulator logs packet counts for all edges and nodes that would not be available in an actual deployment – but that are available in a simulated environment and useful for evaluation purposes. This includes ${s}_{e}$, ${d}_{e}$, ${\rho}_{e}$, and ${\beta}_{e}$ for all edges as well as ${\rho}_{j}$ for all nodes and gateways. We evaluate the proposed reliability estimation protocols by comparing the $\hat{\rho}_{j}$ output by the protocol to the underlying ground truth reliability ${\rho}_{j}$, both in settings where packet drops are due to random failures and settings where packet drops are malicious.
A.3.4 Simulation Runtime
@@
A.3.5 Experimental Setup: Honest but Unreliable Setting
@@
[…]
@@
We set the probability of a packet being a measurement to $1\%$, i.e., $p_{\mathsf{lot}}=0.01$. We vary the total amount of packets encoded by clients to study the node reliability estimation accuracy relative to the number of measurement samples throughout the epoch. The results are shown in Figure 2.
A.3.6 Experimental Setup: Adversarial Setting
@@
To simulate adversarial settings we consider scenarios with $A=\{1,2,4,8,16,32,64\}$ adversaries and $T=\{1,2,4,8,16,32,64\}$ targets. We only consider node assignments to layers that are most favorable to the adversary, i.e., where all $A$ adversarial nodes are in layers adjacent to all the $T$ targets, either as predecessors, as successors, or both. In cases where all $A$ nodes are in a single layer, the $T$ nodes may be all in the preceding layer, all in the succeeding layer or, for $T>1$, distributed as $\frac{T}{2}$ in the preceding and $\frac{T}{2}$ in the succeeding layers. In cases where the $A$ nodes are distributed among two layers, we consider that all $T$ targets are in the layer between the adversaries, with $\frac{A}{2}$ as predecessors and the other $\frac{A}{2}$ as successors of the $T$ targets. For $64$ adversaries or targets we only consider scenarios where they are split over two layers (rather than all in one layer). We consider that the $A$ adversaries maximize the attack impact by dropping all (rather than just some) of the packets in the edges shared with the $T$ targets – while not dropping any packets on edges shared with non-targets. Targets are reliable nodes that do not drop any packets and thus have a true reliability of $100\%$. For the remaining ‘vanilla’ nodes, which are non-adversarial and non-targets, we consider two cases: one where they are all reliable (results reported in Section 4.1.2), and another where $30\%$ of the nodes in each layer are potentially unreliable (results reported in Appendix A.3.7). The second case uses a setup almost identical to the one described in Section A.3.5, with the only difference being that only $16$ (rather than $32$) unreliable nodes per layer may go offline during the epoch. We consider scenarios with a total of $2$ million measurement packets per epoch of one hour.
@@
Figure 4: Each sample represents the aggregate reliability penalties $c_{T}$ for $1\leq|T|\leq 64$ targets ($y$ axis) and $c_{A}$ for $1\leq|A|\leq 64$ adversaries ($x$ axis) in $630$ simulation runs. Purple circles represent scenarios where adversaries control at most $20\%$ ($16$ out of $80$ nodes) of any layer. Green ‘+’ signs correspond to simulations where adversaries make up $40\%$ ($32$ out of $80$ nodes) of at least one layer.
A.3.7 Experimental Results: Settings that Combine Adversarial and Unreliable Nodes
We finally examine scenarios where some nodes not involved in the attack are unreliable, as described in the experimental setup in Appendix A.3.6. Similar to the evaluation in Section 4.1.2, we compute the adversarial cost for $|A|$ malicious nodes as $c_{A}=|A|-\sum_{i\in A}\hat{\rho}_{i}$ and the cost imposed on the $|T|$ targets as $c_{T}=|T|-\sum_{j\in T}\hat{\rho}_{j}$. Each simulation run produces a result sample $(c_{A},c_{T})$ that we represent as a dot in a scatter plot. We represent scenarios with large number of adversaries with a different symbol to better evaluate the impact of high corruption rates. The results are shown in Figure 4, where we can see that adversary and target penalties are still symmetric in the majority of scenarios. We can see that $c_{A}\approx c_{T}$ in all scenarios below a certain scale. As the attack scale (proportional to $|A|$x$|T|$) increases however, there are increased chances that unreliable nodes affect the measured reliability of either adversaries or targets, making the attack cost asymmetric. This happens in particular when the combination of attack-related drops and unreliability affects the median incoming or outgoing link reliability of a node, i.e., when more than half of a node’s incoming or outgoing links have sub-par performance for one or another reason. This makes the node be labeled as unreliable and increases the attribution of drops to that node. This can more easily happen to targets that are under attack from a large number of adversaries in an adjacent layer, as well as to adversaries that attack a large number of targets.
On the left side of Figure 4 above the diagonal we see a set of results for which the adversary succeeds in causing a large reliability degradation on targets at no penalty to itself ($c_{A}\ll c_{T}$). This degree of adversarial success requires controlling $40\%$ ($32$ out of $80$ nodes) of at least one layer (all those samples are green ‘+’ symbols) but also some luck for the adversary. In various other instances with the same degree of adversarial control we observe that either: (i) the adversary has a penalty $c_{A}$ that is lower than that of the targets’ $c_{T}$, but still proportional to it and non-negligible (row of green ‘+’ symbols over the diagonal for which $c_{A}<c_{T}$); (ii) the adversary’s penalty is symmetric with the target’s (‘+’ symbols on the diagonal for which $c_{A}\approx c_{T}$); or (iii) in a few cases the adversary pays a large penalty $c_{A}>10$ without inflicting any penalty on the targets (‘+’ symbols on the bottom of the plot for which $c_{A}\gg c_{T}$).
The considered level of unreliability is not helpful for adversaries that control up to $20\%$ of any layer (purple circle symbols). In these cases, at best the adversary pays a symmetric penalty (results on the diagonal, for which $c_{A}\approx c_{T}$), while in other instances it pays a higher cost than the target (circles below the diagonal for which $c_{A}>c_{T}$). In some instances the adversary pays a large penalty without the attack having an effect on the target (circles on the bottom of the figure for which $c_{A}\gg c_{T}$).